OpenStudy (aimmebgood):

12x^(2) (x + 3) + 7x (x + 3) - 45 (x + 3) factor completely. I got as far as factoring out the (x + 3). where do you go from there?

OpenStudy (anonymous):

\[(x+3)(12x^2+7x-45)\]

OpenStudy (anonymous):

What are you left with after you factor out (x+3)? Or, see above :)

OpenStudy (anonymous):

going to need to use quadratic or maybe you can split up the middle i'd pick quadratic since your a and c are pretty big.

OpenStudy (anonymous):

\[x=\frac{-b\pm \sqrt {b^2-4ac}}{2} \] for \[0=ax^2+bx+c\]

OpenStudy (anonymous):

\[\frac{-7\pm \sqrt{49-4(12)(-45)}}{24}\]

OpenStudy (anonymous):

the denominator of the formulae shoudl be 2a not just 2

OpenStudy (aimmebgood):

I have no idea what that is outkast.

OpenStudy (anonymous):

\[\frac{-7\pm \sqrt{2209}}{24}=\frac{-7\pm 47}{24}\]

OpenStudy (anonymous):

quadratic formulae is a way of finding zeroes/roots. Usually it's used in a sense where you have a complex or rational zeroes/roots.

OpenStudy (aimmebgood):

Im sorry. Im not that smart when it comes to math, so I don't really understand anything that you're explaining.

OpenStudy (anonymous):

when you factor something you are trying to basically find as many roots/zeroes as you can. a zero/root is when you equate a multiple of equation to 0, the whole equation will become 0. For example, before you did anything , you pulled out a (x+3). -3 is a root because if you let x=-3, (x+3)=0, which then would make \[(x+3)(12x^2+7x-45)=0\] \[0(12x^2+7x-45)=0\] \[0=0\]

OpenStudy (anonymous):

what the quadratic formulae does is find the zeroes/roots just like x=-3 so that you can factor it completely. Say if you use the quadratic formula above and you get x=5, this would mean that \[x+c=0\] where c is just a number. Since we know x=5 we now get \[5+c=0\] subtracting 5 would get \[c=-5\] so now we know that \[x-5=0\] when x=5, and we can factor (x-5) out now

OpenStudy (anonymous):

if this is new to you we'll just use the other method so take \[12x^2+7x-45=0\] where \[ax^2+bx-c=0\] multiply a*c \[12*-45=-540\] now find factors of -540 that add up to b, which equals 7